# Request a Note: How the Request Function Shapes X’s Community Notes System

### v1

@@ 4.3. Algorithmic Evaluation of Community Notes
Factors associated with note polarization. The estimation results for a linear regression model predicting note polarization are shown in Fig. 7. Community notes that cite politically unbiased domains tend to have reduced polarization across their ratings (coef.=0.019coef.=-0.019, $p<0.001$), whereas citations to domains with higher information quality are associated with increased polarization ($coef.=0.043$, $p<0.01$).
Regarding post content, positive sentiment in posts is negatively associated with note polarization (coef.=0.046coef.=-0.046, $p<0.001$). Posts related to Politics are linked to higher polarization ($coef.=0.053$, $p<0.001$), whereas posts related to Finance & Business (coef.=0.021coef.=-0.021, $p<0.001$) are associated with lower polarization. Community notes on posts containing stronger claims ($coef.=0.037$, $p<0.001$) or more opinionated content ($coef.=0.063$, $p<0.001$) tend to be more polarized. In contrast, community notes on posts that include media (coef.=0.026coef.=-0.026, $p<0.001$) or exhibit higher misleadingness (coef.=0.030coef.=-0.030, $p<0.001$) have lower polarization scores. For post authors, posts from all types of verified accounts—blue ($coef.=0.044$, $p<0.001$), business ($coef.=0.041$, $p<0.001$), and government ($coef.=0.053$, p<0.001p<0.001)—tend to receive community notes with higher polarization compared to posts from unverified accounts. Additionally, posts authored by accounts with more followers are associated with greater note polarization ($coef.=0.057$, $p<0.001$). However, posts from accounts with higher misinformation exposure scores are more likely to receive community notes with reduced polarization (coef.=0.163coef.=-0.163, $p<0.001$).
In summary, the polarization of community notes is positively associated with (i) high quality of external domains, (ii) political topics, (iii) posts containing claims or opinionated content, and (iv) verified accounts with many followers. Conversely, the polarization of community notes is negatively associated with (i) citations to unbiased domains, (ii) positive sentiment, (iii) posts containing media, (iv) misleadingness of post content and (v) misinformation exposure scores of post authors.
Together with our findings in Section 4.2, which show that requests shift top writers’ attention toward political content while prioritizing posts with higher potential misleadingness, the analysis of factors affecting note helpfulness and polarization provides additional insight. We find that post misleadingness is positively associated with note helpfulness but negatively associated with note polarization, whereas political content exhibits the opposite pattern. Consequently, the higher helpfulness and lower polarization of request-fostered notes can be partially interpreted as reflecting top writers’ selective fact-checking of misleading posts.
@@ 4.4. Summary of Main Findings
  • Requested posts are more likely to receive notes when they are associated with entertainment, finance, media, high misleadingness, or authors with high misinformation exposure, but less so for political or opinionated content. This suggests that Community Notes contributors and requestors have distinct selection patterns.
@@ 5.1. Research Implications
Content selection by contributors. Our results show that contributors’ decisions to fact-check requested posts are shaped by both post content and author characteristics. Specifically, contributors are less likely to write notes for posts with negative sentiment or opinionated content, authored by blue or business accounts, or focused on politics or science. In contrast, posts related to entertainment or finance, those containing claims or media, and those flagged by GPT as more misleading are significantly more likely to receive notes. At the post author level, posts from accounts with higher misinformation exposure or right-leaning partisanship are also more likely to be annotated. [^1] To validate our findings and rule out confounding from request-driven activity, we exclude posts with request-fostered community notes (as identified in Section 4.2) and repeat our analysis. The results remain robust and consistent (see details in Suppl. S2). Notably, in this robustness check, the presence of claims becomes positively associated with the likelihood of receiving a community note, further reinforcing the role of substantive content in shaping contributors’ fact-checking decisions. The misalignment between the types of posts surfaced by requestors (e. g., political content) and those contributors prioritize for annotation underscores a divergence in fact-checking selection.
@@
On the one hand, Community Notes contributors tend to fact-check claim posts with higher estimated misleadingness and from authors with higher misinformation exposure scores. This suggests that, compared to requestors, contributors adopt stricter standards, prioritizing content with greater potential for misinformation and harm. On the other hand, previous research finds that contributors often fact-check posts that are relatively straightforward to verify or contain claims already addressed in earlier expert fact-checks (Borenstein et al., 2025). Similarly, LLMs perform strongly when processing logically structured claims but struggle with more complex ones (Costabile et al., 2025). This suggests that, although contributors appear to focus their attention on posts with higher GPT-estimated misleadingness, their fact-checking may simultaneously be biased toward content that is easier to assess. Such a focus can improve efficiency and support the production of high-quality notes, but it also risks overlooking politically salient or ambiguous posts that requestors perceive as urgent. Addressing this gap may require hybrid models of collaboration, where professional fact-checkers complement contributors by targeting content that is more ambiguous, emergent, or difficult to verify (Augenstein et al., 2025).
@@
Partisan selection biases between contributors and requestors. Notably, the human-centered fact-checking approaches—whether expert-based or crowdsourced—cannot fully avoid political bias in the selection of targets (Chuai et al., 2025b; Pilarski et al., 2024; Shin and Thorson, 2017). Our results reveal significant differences in how requestors and contributors approach political content and partisan cues. Previous work has shown that users preferentially challenge content authored by those with opposite partisan leanings, and Republicans are flagged more often than Democrats for sharing misinformation on $\mathbb{X}$’s Community Notes (Renault et al., 2025; Kuuse et al., 2025; Allen et al., 2022). Our findings echo this pattern: contributors are more likely than requestors to write notes on posts authored by Republicans, whereas requestors are comparatively more likely to surface posts authored by Democrats. This divergence raises an important interpretive challenge. On the one hand, it may reflect partisan selection biases, with anonymized contributors and requestors each applying their own subjective judgments when deciding what to fact-check (Martel et al., 2025). On the other hand, it could mirror underlying asymmetries in the misinformation landscape, where different political groups are disproportionately represented in the spread of misleading content. The coexistence of such tendencies highlights the complexity of maintaining balance and fairness in decentralized fact-checking systems like Community Notes.
@@
Reliability and vulnerability of the note selection algorithm. We provide a comprehensive analysis of how the helpfulness and polarization of community notes are shaped by both content and source factors. Posts with higher estimated misleadingness and authored by accounts with greater misinformation exposure are more likely to elicit helpful annotations. This finding supports the view that Community Notes can function as a reliable mechanism for surfacing accurate fact-checks and correcting misinformation (Allen et al., 2024).
@@ 5.3. Limitations and Future Work
Our study has several limitations that future research could address. First, our analysis is based on observational data from $\mathbb{X}$, which constrains causal inferences. While we examine various characteristics—related to posts, their authors, and associated notes—and identify their associations with note outcomes, we cannot definitively determine whether certain factors directly cause higher helpfulness or lower polarization. Second, while GPT-based measures provide scalable assessments of potential misinformation, they may not perfectly address emerging and complex misleading content. This could affect interpretations regarding contributors’ priority strategies. Third, given that the top-writer status is not available in our dataset, we cannot perfectly distinguish request-fostered notes from contributor-driven notes. Nevertheless, our estimation method ensures that the effect of the request feature in fostering note writing is not underestimated. Finally, contributors and requestors are anonymized in the Community Notes system, preventing the analysis of individual-level behavior, motivation, or expertise. While this reduces rater bias, it also limits understanding of how personal experience or identity may influence content selection and note quality.
@@ Acknowledgements.
This research is supported by the Luxembourg National Research Fund (FNR) and Belgian National Fund for Scientific Research (FNRS), as part of the project REgulatory Solutions to MitigatE DISinformation (REMEDIS), grant ref. INTER_FNRS_21_16554939_REMEDIS.
Supplementary Materials for
“Request a Note: How the Request Function Shapes X’s Community Notes System”
Appendix S1 Evaluation of GPT Outputs
Previous work has shown that Large Language Models (LLMs) can effectively identify claims and perform well in truthfulness classification, underscoring their potential to support crowdsourced fact-checking (Costabile et al., 2025; Choi and Ferrara, 2024; DeVerna et al., 2024). Building on this, we employ GPT-4.1, the state-of-the-art GPT model at the time of this study, to assess whether posts contain factual claims and/or personal opinions, and to estimate the misleadingness of each post. The GPT outputs are based on the following prompt:
GPT prompt for content type and misleadingness You are an assistant that evaluates tweet content for two independent dimensions: content type and misleadingness. For each input tweet, return a single-line JSON string with the following structure: { “type_scores”: { “claim”: float (0 to 1), # Degree to which the tweet presents a factual/verifiable claim “opinion”: float (0 to 1) # Degree to which the tweet expresses a subjective opinion }, “misleadingness”: float (0 to 1) # How misleading the tweet is, where 0 = not misleading, 1 = extremely misleading } Guidelines: - “claim” and “opinion” scores are independent and do not need to sum to 1. - Only return the JSON string. Do not include any explanation or additional text.
Notably, GPT-generated annotations are not treated as definitive ground truth. Instead, they function as standardized proxies that allow for large-scale comparisons across posts. We manually validate a random subsample of $200$ posts and evaluate the performance of the GPT outputs. Additionally, to balance workload, six trained research assistants each independently evaluate a set of $100$ posts, with each post reviewed by three assistants. Specifically, for claims and opinions, assistants rate the extent to which a tweet presents a factual or objectively verifiable claim or expresses a subjective opinion, using a scale from 0 to 1. For misleadingness, assistants rate the accuracy of the GPT-generated explanation on the same 0–1 scale. The explanation is generated by GPT based on the following prompt:
GPT prompt for misleadingness explanation You are an assistant that previously predicted the misleadingness of a given tweet, returning a score between 0 and 1. Your task now is to provide a concise and factual explanation that justifies the specific misleadingness score you gave to the tweet. Do not re-score the tweet. The explanation must be no more than 100 words.
We examine the Pearson correlation coefficients between the scores assigned by research assistants and those generated by GPT for claims and opinions. For each post, we average the ratings from the research assistants and compute the correlation with the corresponding GPT scores. The correlations are 0.403 ($p<0.001$) for claims and 0.620 ($p<0.001$) for opinions, indicating a substantial alignment between human judgments and GPT outputs.
To further assess reliability, we transform the scores into binary classes using a threshold of 0.5. Specifically, a post is labeled as containing a claim on the human side if at least two of the three assistants assign it a score $\geq 0.5$. Using this criterion, GPT achieves an accuracy of 63% in classifying claims and 73.5% in classifying opinions. In addition, our manual check shows that GPT can provide accurate misleadingness explanations for 90% posts.
Appendix S2 Likelihood of Receiving Community Notes

### v2

@@ 4.3. Algorithmic Evaluation of Community Notes
Factors associated with note polarization. The estimation results for a linear regression model predicting note polarization are shown in Fig. 7. Community notes that cite politically unbiased domains tend to have reduced polarization across their ratings ($\mathit{coef.}=-0.019$, $p<0.001$), whereas citations to domains with higher information quality are associated with increased polarization ($\mathit{coef.}=0.046$, $p<0.01$).
In terms of post content, positive sentiment in posts is negatively associated with note polarization ($\mathit{coef.}=-0.038$, $p<0.001$), but negative sentiment shows the opposite pattern ($\mathit{coef.}=0.017$, $p<0.05$). Posts related to Politics are linked to higher polarization ($\mathit{coef.}=0.054$, $p<0.001$), whereas posts related to Finance & Business ($\mathit{coef.}=-0.022$, $p<0.001$) are associated with lower polarization. Community notes on posts containing stronger claims ($\mathit{coef.}=0.025$, $p<0.001$) or more opinionated content ($\mathit{coef.}=0.041$, $p<0.001$) tend to be more polarized. In contrast, community notes on posts that include media ($\mathit{coef.}=-0.027$, $p<0.001$) or exhibit higher misleadingness ($\mathit{coef.}=-0.032$, $p<0.001$) have lower polarization scores. For post authors, posts from all types of verified accounts—blue ($\mathit{coef.}=0.044$, $p<0.001$), business ($\mathit{coef.}=0.038$, $p<0.001$), and government ($\mathit{coef.}=0.054$, p<0.001p<0.001)—tend to receive community notes with higher polarization compared to posts from unverified accounts. Additionally, posts authored by accounts with more followers are associated with greater note polarization ($\mathit{coef.}=0.057$, $p<0.001$), and posts from accounts with higher misinformation exposure scores are more likely to receive community notes with reduced polarization ($\mathit{coef.}=-0.160$, $p<0.001$).
In summary, the polarization of community notes is positively associated with (i) high quality of external domains, (ii) negative sentiment, (iii) political topics, (iv) posts containing claims or opinionated content, and (v) verified accounts with many followers. Conversely, the polarization of community notes is negatively associated with (i) citations to unbiased domains, (ii) positive sentiment, (iii) posts containing media, (iv) misleadingness of post content and (v) misinformation exposure scores of post authors.
Notably, through the analysis of both note helpfulness and polarization, we observe that post misleadingness is positively associated with note helpfulness but negatively associated with note polarization, whereas political content exhibits the opposite pattern. Together with our findings in Section 4.2, which show that requests shift top writers’ attention toward political content while prioritizing posts with higher potential misleadingness, our analysis provides additional insight: the higher helpfulness and lower polarization of request-fostered notes can be partially interpreted as reflecting top writers’ selective fact-checking of misleading posts.
@@ 4.4. Summary of Main Findings
  • Requested posts are more likely to receive notes when they are associated with entertainment, finance, media, claims, high misleadingness, or authors with high misinformation exposure, but less so for negative, political or scientific content. This suggests that Community Notes contributors and requestors have distinct selection patterns.
@@ 5.1. Research Implications
Likelihood of receiving community notes. Our results show that the likelihood of receiving community notes for requested posts are shaped by both post content and author characteristics. Specifically, posts with negative sentiment, authored by blue or business accounts, or focused on politics or science are less likely to receive community notes. In contrast, posts related to entertainment or finance, those containing claims or media, and those flagged by GPT as more misleading are significantly more likely to receive community notes. At the post author level, posts from accounts with higher misinformation exposure or right-leaning partisanship are also more likely to be annotated. The misalignment between the types of posts surfaced by requestors (e. g., political content) and those annotated by contributors potentially underscores a divergence in fact-checking selection. [^5] To validate our findings on content selection by contributors and rule out confounding from request-driven activity, we exclude posts with request-fostered community notes (as identified in Section 4.2) and repeat our analysis. The results remain robust (see details in Suppl. S6).
@@
On the one hand, claim-oriented posts with higher estimated misleadingness and from authors with higher misinformation exposure scores are more likely to receive community notes. This suggests that, compared to requestors, contributors might adopt stricter standards, prioritizing content with greater potential for misinformation and harm. On the other hand, previous research finds that contributors often fact-check posts that are relatively straightforward to verify or contain claims already addressed in earlier expert fact-checks (Borenstein et al., 2025). Similarly, LLMs perform strongly when processing logically structured claims but struggle with more complex ones (Costabile et al., 2025). This suggests that, although contributors appear to focus their attention on posts with higher GPT-estimated misleadingness, their fact-checking may simultaneously be biased toward content that is easier to assess. Such a focus can improve efficiency and support the production of high-quality notes, but it also risks overlooking politically salient or ambiguous posts that requestors perceive as urgent. Addressing this gap may require hybrid models of collaboration, where professional fact-checkers complement contributors by targeting content that is more ambiguous, emergent, or difficult to verify (Augenstein et al., 2025).
@@
Partisan asymmetries between requests and community notes. Notably, the human-centered fact-checking approaches—whether expert-based or crowdsourced—cannot fully avoid political bias in the selection of targets (Chuai et al., 2025b; Pilarski et al., 2024; Shin and Thorson, 2017). Our results reveal significant partisan asymmetries between reuqested posts and those that ultimately receive community notes. Previous work has shown that users preferentially challenge content authored by those with opposite partisan leanings, and Republicans are flagged more often than Democrats for sharing misinformation on $\mathbb{X}$’s Community Notes (Renault et al., 2025; Kuuse et al., 2025; Allen et al., 2022). Our findings echo this pattern: among requested posts, those authored by Republicans are more likely to receive community notes than those authored by Democrats. This asymmetry raises an important interpretive challenge. On the one hand, it may reflect partisan selection biases, with anonymized contributors and requestors each applying their own subjective judgments when deciding what to fact-check (Martel et al., 2025). On the other hand, it could mirror underlying asymmetries in the misinformation landscape, where different political groups are disproportionately represented in the spread of misleading content. The coexistence of such tendencies highlights the complexity of maintaining balance and fairness in decentralized fact-checking systems like Community Notes program.
@@
Reliability and vulnerability of the note selection algorithm. We provide a comprehensive analysis of how the helpfulness and polarization of community notes are shaped by both content and source factors. Posts with higher estimated misleadingness and authored by accounts with greater misinformation exposure are more likely to elicit helpful annotations. This finding supports the view that Community Notes can function as a reliable mechanism for surfacing accurate fact-checks and correcting misinformation (Allen et al., 2024b).
@@ 5.3. Limitations and Future Work
Our study has several limitations and offers directions for future research. First, our analysis is based on observational data from $\mathbb{X}$, which constrains causal inferences. Although we examine various characteristics—related to posts, their authors, and associated notes—and identify their associations with note outcomes, we cannot definitively determine whether certain factors directly influence contributors’ decisions to write notes and causally lead to higher helpfulness or lower polarization in the resulting notes. For instance, in our post-level analysis (Fig. 2), latent factors—such as unobserved contributor-level confounding and the inherent difficulty of fact-checking a post—are not directly measured or controlled for. These factors may influence both contributors’ selection behavior and the quality of resulting notes. Future studies could explore causal relationships and clarify underlying mechanisms through controlled survey experiments or platform-supported interventions. Second, while our GPT-based approach provides scalable assessments of content types and potential misinformation, it remains subject to hallucinations and model-specific biases (Augenstein et al., 2024), and may not fully capture emerging or highly complex misleading content (Costabile et al., 2025). To mitigate these concerns, we conduct comprehensive validations using both assistant-annotated and expert-annotated datasets to demonstrate the reliability of GPT-generated outputs for our analysis. With continued advances in LLM-based fact-checking approaches, annotation accuracy is expected to further improve. Future work is encouraged to validate our findings when more advanced models are available. Third, given that the top-writer status is not available in our dataset, we cannot perfectly distinguish request-fostered notes from contributor-driven notes. Nevertheless, our estimation method ensures that the effect of the request feature in fostering note writing is not underestimated. Finally, contributors and requestors are anonymized in the Community Notes system, preventing the analysis of individual-level behavior, motivation, or expertise. While this anonymization can reduce rater bias, it also limits understanding of how personal experience or identity may influence content selection and note quality. Future work could explore these dynamics in contexts where such information can be collected ethically through survey experiments or accessed under appropriate privacy safeguards on platforms.
@@ Acknowledgements.
This research is supported by the Luxembourg National Research Fund (FNR) and Belgian National Fund for Scientific Research (FNRS), as part of the project REgulatory Solutions to MitigatE DISinformation, grant ref. INTER_FNRS_21_16554939_REMEDIS.
Supplementary Materials
Appendix S1 Misleading Criteria
The criteria for labeling misleading posts in $\mathbb{X}$’s Community Notes system is shown in Fig. S1. The listed criteria suggest that the Community Notes system aims to provide corrective or contextual information for not only factually incorrect content or alter images and videos, but also posts that present accurate information in a misleading manner, such as, through omission of context, exaggeration, or distortion of the original meaning. This is consistent with the definition of misinformation in our study. Therefore, we consider misleading posts annotated by community notes on $\mathbb{X}$ to be manifestations of misinformation.
Figure S1. The note-writing interface for Community Notes contributors, showing the specific criteria for labeling posts as misleading. Contributors are also required to provide explanations and cite external sources to justify their assessments.
Appendix S2 GPT Prompt
The full GPT-4.1 prompt used in our analysis is provided below. In addition to the definition of misinformation, we adopt the definitions of claim and opinion in a previous study (Ni et al., 2024).
You are an assistant that evaluates tweet content for two independent dimensions: content type and misleadingness.
## Dimension 1: Content Type
Assess whether the tweet expresses a factual/verifiable claim and/or a subjective opinion.
### Claim
A factual claim is a statement that explicitly presents some verifiable facts. Statements with subjective components like opinions can also be factual claims if they explicitly present objectively verifiable facts.
Indicators that a tweet contains a factual claim:
- Mentioning somebody (including the speaker) did or is doing something specific and objective
- Quoting quantities, statistics, and data
- Claiming a correlation or causation
- Assertion of existing laws or rules of operation
- Pledging a specific future plan or making specific predictions about future
### Opinion
An opinion is a judgment based on facts, an attempt to draw a reasonable conclusion from factual evidence. While the underlying facts can be verified, the derived opinion remains subjective and is not universally verifiable.
## Dimension 2: Misleadingness
Assess how misleading the tweet is, i.e., its potential to express misinformation, regardless of intent.
###Definition
Misinformation refers to false or misleading information that is contradicted by empirical evidence or inconsistent with common or expert consensus, without imposing assumptions on the intent or the format of the content
Indicators of misinformation or misleadingness:
- It contains a factual error
- It contains a digitally altered photo or video
- It contains outdated information that may be misleading
- It is a misrepresentation or missing important context
- It presents an unverified claim as a fact
- It is a joke or staire that might be misinterpreted as a fact
## Output Format
For each input tweet, return a single-line JSON string with the following structure:
“type_scores”: {
“claim”: float (0 to 1), # Degree to which the tweet presents a factual/verifiable claim
“opinion”: float (0 to 1) # Degree to which the tweet expresses a subjective opinion
“misleadingness”: float (0 to 1) # How misleading the tweet is, where 0 = not misleading, 1 = extremely misleading
## Guidelines:
- ”claim” and ”opinion” scores are independent and do not need to sum to 1.
- Only return the JSON string. Do not include any explanation or additional text.
Appendix S3 GPT Validation Using PolitiFact
In addition to the manual validation conducted with student assistants, we further evaluate the performance of our GPT-based approach using expert-verified data from PolitiFact. Specifically, we collect all statements fact-checked by PolitiFact between July 15, 2024 and June 4, 2025, aligning with the time window of our dataset of requested posts. We obtain $1320$ statements in total and retain those rated as (mostly) false (== $1016$) or (mostly) true (== $49$), resulting in $1065$ statements for evaluation. Applying the same GPT-based annotation pipeline, we compare GPT assessments with PolitiFact’s expert verdicts. GPT achieves a high performance ($\mathit{Accuracy}=$ 0.801; $\mathit{Weighted~F1}=$ 0.857), consistent with the results of our manual validation. Notably, despite the substantial class imbalance toward false statements, GPT attains comparable recall for true ($\mathit{Recall}=0.816$) and false statements ($\mathit{Recall}=0.800$), demonstrating its effectiveness in identifying both accurate and misleading content.
Appendix S4 Likelihood of Receiving Community Notes for Political Posts
Fig. S2 presents the estimation results for political post regarding the likelihood of receiving community notes. The results remain robust and consistent with our main findings.
Appendix S5 Robustness Check for Top Writers’ Post Selection
In the main paper, we find that top writers are more likely to select political posts and posts with higher potential misleadingness after the request threshold, compared to before the request threshold. To ensure the robustness of our findings and reduce potential unobserved confounding related to potential heterogeneity among top writers, we restrict the analysis to posts selected by the same set of top writers before and after the request threshold. The estimation results are presented in Fig. S3. The estimates for political posts ($\mathit{coef.}=0.162$, $\mathit{odds~ratio}=1.176$, $p<0.001$) and misleadingness ($\mathit{coef.}=0.413$, $\mathit{odds~ratio}=1.511$, $p<0.001$) remain consistent with our findings, further supporting that requests shift contributors’ attention toward political content and posts with higher misleadingness.
Appendix S6 Robustness Check for Likelihood of Receiving Community Notes
@@
Figure S2. The estimation results for the logistic regression model predicting the likelihood that a requested political post receives a community note. Shown are coefficient estimates with 95% CIs. The coefficients for misinformation exposure score and partisan score are estimated based on the subset of posts containing the corresponding information. The number of words in each post is controlled during estimation but omitted in visualization for better readability. Continuous independent variables—word count, account age, number of followers, and number of followees—are z-standardized before estimation to facilitate interpretation.
Figure S3. The estimation results for the logistic regression model predicting the likelihood that a requested post receives a top-writer note after the request threshold, i. e., request-fostered notes. Shown are coefficient estimates with 95% CIs. The coefficients for misinformation exposure score and partisan score are estimated based on the subset of posts containing the corresponding information. The number of words in each post is controlled during estimation but omitted in visualization for better readability. Continuous independent variables—word count, account age, number of followers, and number of followees—are z-standardized before estimation to facilitate interpretation.